a particle is project ed from ground with initial velocity of 20 m/s  at an angle of 30 degree with horizontal.the magnitude of change in velocity in time interval from  t=0 to t=0.5 sec is

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

only vertical component of velocity changes.  Initial value of vertical component =10m/s upward. when t=0.5s, vertical velocity=10-(g*0.5)=5m/s upwards,magnitude of change in velocity=5m/s